Find the magnitude of the charge if it has a Voltage of 11 V with a distance of 23 mm.

1
Expert's answer
2022-05-12T15:09:34-0400
"V=k\frac{q}{r}"

"q=\frac{Vr}{k}=\frac{11*0.023}{9*10^9}=2.8*10^{-11}\:\rm C"